# StreakNet-Arch: An Anti-scattering Network-based Architecture for Underwater Carrier LiDAR-Radar Imaging

### v2

@@ IV-E DBC-Attention is more suitable for underwater optical 3D imaging than Self-Attention
To demonstrate the superiority of DBC-Attention over Self-Attention in underwater imaging tasks, we conducted ablation experiments by replacing the Self-Attention module in StreakNet with DBC-Attention while keeping all other parameters unchanged. Through the experimental results (Fig. 7 and Table III), we found:
@@
We simultaneously conducted ablation experiments with traditional imaging methods. We performed Attention Analysis on the FD Embedding Layer of both StreakNet and StreakNetv2. The results were used as equivalent filters, replacing the traditional 450 MHz - 550 MHz bandpass filter for imaging. The results on the validation set are presented in Fig. 14b and Table V.
@@
Model Traditional (baseline) StreakNet-Emb. StreakNetv2-Emb.
@@
V Conclusion
This paper tackles two significant challenges impeding conventional underwater imaging algorithms: limited anti-scattering capabilities and the inability to achieve real-time processing. We incorporated Self-Attention mechanism networks into the signal processing stage to enhance the scatter resistance capability of the UCLR. Primarily, we introduce StreakNet-Arch, a groundbreaking end-to-end binary classification architecture, to overcome the real-time imaging hurdle. Additionally, DBC-Attention, a novel variant of the Self-Attention mechanism specifically optimized for underwater environments, is proposed. Experimental results conclusively demonstrate DBC-Attention’s superiority, achieving significantly better performance compared to the standard Self-Attention approach. Furthermore, we have released the first-ever public dataset containing 2,695,168 real-world underwater 3D point cloud data of streak-tube camera images. This work paves the way for future research on optimizing filtering methods in underwater imaging algorithms.

### v3

@@ IV-E DBC-Attention is more suitable for underwater optical 3D imaging than Self-Attention
To demonstrate the superiority of DBC-Attention over Self-Attention in underwater imaging tasks, we conducted ablation experiments by replacing the Self-Attention module in StreakNet with DBC-Attention while keeping all other parameters unchanged. Through the experimental results Table IV), we found:
Figure 13: Overview of the field experiment site. (a) Field Experiment Location (Google Maps view). (b) Bathymetry of the Field Experiment Area.
@@
We simultaneously conducted ablation experiments with traditional imaging methods. We performed Attention Analysis on the FD Embedding Layer of both StreakNet and StreakNetv2 (Use StreakNet-Emb. and StreakNetv2-Emb. to denote them, respectively). The results were used as equivalent filters, replacing the traditional 450 MHz - 550 MHz bandpass filter for imaging. The results on the validation set are presented in Fig. 11b and Table V.
@@
Model Bandpass (baseline) StreakNet-Emb. StreakNetv2-Emb.
@@
V Field Experiment
To evaluate the imaging performance of the UCLR system in deep-sea conditions, a field experiment was conducted on October 29, 2023, aboard the Dongfang Haike research vessel in the South China Sea (E 110° 12.123’, N 17° 20.521’, Fig. 13a). The bathymetry at the experimental site is approximately 1200 meters (Fig. 13b). During the experiment, conducted under Sea State 3 (slight seas, $\leq$ 1.25 m waves), the prototype system was deployed to a depth of 1000 m using a ship-mounted winch, and the target was suspended 20 m beneath it via an iron chain (Fig. 14a, b). As shown in Fig. 14e, the target measures 1000 mm $\times$ 1000 mm, with a 400 mm $\times$ 400 mm raised platform of 600 mm height at the center.
Figure 14: Setup and Results of the Field Experiment. (a) Schematic of the field experiment setup and the target object. (b) On-site photo taken during the field experiment. (c) 3D imaging results obtained from the 1000 m deep-sea experiment. (d) The measured height of the protruding platform is 554 mm, with an absolute error of 46 mm compared to the ground-truth value, corresponding to a relative error of 7.6%. (e) Schematic of the target object.
Due to the challenge of manually calibrating scatter suppression for binary classification ground truth in deep-sea environments, we utilized the StreakNetv2-m model, previously trained with water tank data, to perform 3D imaging. The relative error between the measured and true height of the target protruding platform after imaging was used as the evaluation metric for imaging performance.
After imaging with the UCLR system, the measured height of the protruding platform was 554 mm (Fig. 14c, d), with an error of 46 mm compared to the true value, resulting in a relative error of 7.6%. The results of the field experiment validate the applicability of the UCLR system in deep-sea environments.
VI Discussion
Although StreakNet-Arch, particularly StreakNetv2 based on DWC-Attention, demonstrates superior imaging quality in the water tank environment compared to traditional Bandpass filtering, MP models, CNNs, and Self-Attention-based StreakNet within a certain computational complexity range, the StreakNetv2 network still presents a notable risk of overfitting. A contributing factor is that the current training set consists of high-resolution 3-D point clouds that must be painstakingly hand-labeled, leaving the model dependent on fully supervised learning. For example, StreakNetv2-l and StreakNetv2-x, when reaching a computational complexity of 10 GFLOPs (Table II), achieve lower F1 scores than the smaller MP and CNN models (Table IV). Nevertheless, those human-annotated labels enable the network to learn far richer spatial–temporal correlations than traditional algorithms can capture: supervised StreakNet-v2 not only delivers markedly higher imaging fidelity but also sustains real-time throughput, thereby retaining a decisive edge in both quality and speed. These results motivate future work on unsupervised or self-supervised formulations that can alleviate the annotation burden while preserving, or even enhancing these performance gains.
VII Conclusion
This study addresses two longstanding bottlenecks in underwater imaging, pronounced susceptibility to scattering and limited real‑time throughput, by embedding self‑attention mechanisms directly into the self-developed UCLR’s signal‑processing pipeline. Building on this integration, we present StreakNet‑Arch, an end‑to‑end binary‑classification framework, and DBC‑Attention, a bespoke self‑attention variant optimized for turbid aquatic scenes. Together, these innovations markedly enhance scatter resistance while sustaining real‑time performance, thereby establishing a new benchmark for high‑speed, high‑fidelity underwater imaging.
Extensive experiments on our validation set under controlled water tank environment demonstrate that both the Self-Attention-based StreakNet and the DBC-Attention-based StreakNetv2 substantially outperform traditional bandpass filtering, and achieve higher $F_{1}$ scores than learning-based MP networks and various CNN models with comparable model sizes and computational complexity. In real-time benchmarks on an NVIDIA RTX 3060 GPU, the proposed StreakNet-Arch maintains a constant Average Imaging Time (AIT) of 54 to 84 ms regardless of the number of input frames, whereas traditional algorithms’ AIT grows linearly, from 58 ms at $N=2$ to 1,257 ms at $N=64$, confirming StreakNet-Arch’s clear advantage for large-scale, real-time imaging.
To foster further progress, we release the first public dataset of 2,695,168 real-world underwater 3D point clouds captured by streak-tube camera. Finally, we validate the complete UCLR system in a deep-sea trial in the South China Sea, achieving an error of 46 mm at 1,000 m depth and 20 m target range. This work not only sets new benchmarks in anti-scattering performance and real-time throughput but also provides a foundation for future advances in underwater imaging filtering strategies.